OpenStudy (anonymous):

If (2x-2)(2-x)=0 , what are all the possible values of x? A. 0 only B. 1 only C. 2 only D 1 and 2 only E. 0, 1 and 2

OpenStudy (anonymous):

2x^2 - 2x+4=0 and Discriminant ( x= -1 or x = 4 ) ???

OpenStudy (whpalmer4):

Okay, if you've got a product of two things that equals 0, you know that one or both of them must also equal 0, because multiplying by 0 is the only way to get 0 in a multiplication. (2x-2) = 0 2x = 2 x = 1 (2-x) = 0 2 = x So x can have the values 1 and 2 only to make this equation true.

OpenStudy (whpalmer4):

0 is included in your answer list as a red herring. We can that it does not satisfy the equation: (2(0)-2)(2-0) = 0 (-2)(2) = 0 -4 = 0 Nope!

OpenStudy (anonymous):

:) yeees thank you very much :)))

OpenStudy (whpalmer4):

\[(2x-2)(2-x) = 2x(2-x) - 2(2-x) = 4x - 2x^2 -4 -(-2x) =\]\[ -2x^2 + 8x - 4\]

OpenStudy (anonymous):

oops yeah :))

OpenStudy (whpalmer4):

Everyone makes mistakes. Well, I don't. I never make misteaks ;-)

OpenStudy (anonymous):

hahah good for you :)))